OpenStudy (anonymous):

Distance = Rate(Time) or d = r t. Find d if r = 65 mph and t = 7 hrs.

OpenStudy (anonymous):

455miles

OpenStudy (anonymous):

\(\Huge \frac{65miles}{\cancel{hour}} \cdot 7 \cancel{hours}=(65 \cdot 7) miles\)